Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Fetzer, Bobbie Meng, Hilton Soto, Geoffrey Stewart Announcements
- Discussion of the tech job market.
- Campus Network Security new level policy on DDD list outlining security it was sent to a review committee they had it 3 hours, then it was sent to an approval committee. Net Managers List sent out another version, complete with super-secret details. This security issue is top priority. There was a four-page list of CLAST machines that were in violation. Geoffrey had the misfortune of powering up a machine to patch just when they were running a scan and it was caught. There is talk of dismissing the department heads in the departments that are not in compliance with the security policy. ITSA Day Bill recommends attendance. Security Training (hosted by CNS & UF Personnel) Bill will advise all liaisons to attend.
- Gators beat Kentucky. Jags lost.
Ongoing Systems Updates
- Status of the West soil sample is still unknown. Shorb did not show up at the last Update Meeting. BDF storage space no plans, they are just going to start. PPD memo was substituted for plans, even though the memo showed up 2 days after the plans were due. Networking should be planned; Systems will require a viable floor plan.
- People that moved from West to East Accountings furniture did not fit. Three electric erasers were found in the obsolete equipment - Suzy was able to identify. Laydene found a memo circa 1935.
- Cell phones issued in Systems. Directories available to us only. Belt holders have been ordered. Discussion of transferring calls to cells: Link + 9 + #. Desk phones may disappear without notice.
- East south tower reconfiguration fixed.
- 4th floor wiring - no encumbrance received by CNS yet. This needs to happen right after this meeting.
- North tower reconfiguration nothing final.
- Reader printers from Microform to MSL held up; maybe 6 will go.
Software Projects Status
- Orders nothing major.
- Need to order large screens for PeopleSoft users (Business Services) will still require scrolling, no matter what. We will also include the extreme programmers in that order.
- Hardware AC in server room for the security servers have not heard anything.
- Geoffrey received a returned Toshiba Techtra from John Ashcraft. He will get it up to par for check out and send the specs to Suzy.
- Storage Retrieval Patron side is working, enhancements ongoing to staff side.
- Lassie licensing Docutek no word.
- Maps/GIS on hold until Yings return.
- ALEPH holding.
- OPS Tracker no news.
Student and Department Happenings
- Mary Beth Suzy reports that the committee has not yet written a recommendation, but she will get one. She also has a traveling spouse, so there are other factors to negotiate.
- Keep a close eye on the physical condition of the building. If anything needs fixing, stay on top of it get requests to Facilities immediately.
Back to the Systems Homepage
Last updated October 7, 2003
by Debra Fetzer